OpenStudy (anonymous):

is anbody good in geometry and is willing to help me out? perferably people who go to connexus

OpenStudy (nikato):

I don't go there, but if you post the question, I can try to help or other people who will see it

OpenStudy (anonymous):

The figure below are similar. What are a) the ratio of the perimeters and b) the ratio of the areas of the larger figure to the smaller figure? The figures are not drawn to scale

OpenStudy (anonymous):

u belive in god

OpenStudy (nikato):

Was that all that was given? What about the other sides?

OpenStudy (nikato):

Nvm you don't need that. The ratio of the perimeter of the small to the big is just 6:26 which can be simplified to 3:13 If you need big to small, just switch them around so 13:3

OpenStudy (anonymous):

yes I do and okay thank you! what about b?

OpenStudy (nikato):

For b, Just square the ratio 13:3 And you get 169:9 for ratio of area

OpenStudy (anonymous):

what about The area of a regular hexagon is 40 cm^2. What is the area of a regular hexagon with sides five times as long?

OpenStudy (anonymous):

Ration large side/small side = 5 Ratio: area L/area S = 25 Area of the larger hexagon: 40 x 25 = 1,000 cm^2

OpenStudy (anonymous):

thank you! okay um a few more. What is the area of a sector with a central angle of 210 degrees and a diameter of 4.6 m?

OpenStudy (anonymous):

Area of the circle : D^2*Pi/4 = Pi* (4.6)^2/4) = 5.29*Pi m^2 Area of the sector is equal to (210/360 = 7/12) of the circle area Area = 5,29*P1*(7/12) = 3.09*Pi m^2